# WLM: Dynamics of an isolated Dwarf Irregular Galaxy Under Ram Pressure in the Local Group

### v1

@@ 3 HI distribution in WLM
A visual inspection of our high resolution moment-0 map in Fig. 2 shows us that WLM has a dense central structure which tapers into the rest of the gaseous disk. This is also seen in the corresponding PV diagram in Fig. 3, where the dense gas core is seen along the steep velocity gradient at the center of the galaxy.
@@
Higgs et al. (2021) found that the stellar radial profile of WLM in their g- and i-band survey is better fitted by 2 distinct morphological components instead of a single one, both components being described by a compact Sersic profile in the inner region and a shallower Sersic profile in the outer region. They tested whether a single component describes the profile better than two components by using the Akaike Information Criteria and the Baysian Information Criteria. In their study for WLM a two-component model is almost 600 times more likely than the single component model, significantly favoring the two-component case. In the following, we repeat this test for the 1D gas profile of WLM.
@@
Classical ellipse fitting to obtain a 1D HI flux profile may not work well for a galaxy at a very high inclination such as WLM. Therefore, we use two separate methods to extract the 1D profile; We perform the ELLPROF task in BBarolo (Teodoro and Fraternali 2015), which uses standard tilted-ring models, and we also implement the following simple procedure to get a profile to test for consistency.
@@
The semi-major axis of the galaxy is determined by measuring the flux in the moment-0 map along a straight line along the PA reported in Table 1. The line is terminated when the flux reaches the RMS noise in the moment map. The RMS noise in the map is determined by the noise in one channel multiplied by the square root of the number of contributing channels (Ianjamasimanana et al. 2018).
@@
The semi-minor axis is calculated with the same procedure repeated along a line perpendicular to the PA along the center of the galaxy and thus an axis ratio is calculated. A central pixel is picked, corresponding to the chosen dynamical center, which is derived in Section 4.3, and ellipses are placed on the image, their size starting with the axes lengths and then decreasing with a set value, while maintaining the axis ratio and PA of each. The image is broken into annular regions, with each ellipse being an annular boundary, mean flux is calculated for each region, so is the standard error on the mean value. Note that the ellipses are not being fit, they are determining spatial bin regions.
@@
In equation 1, SE, and n are the standard error, the standard deviation of the flux values and the number of pixels within each in the annulus, respectively.
@@
Thus total uncertainty is calculated for each annulus with Eq. 2, with the RMS being calculated as described earlier in this section.
@@
These values are then corrected for inclination presented in Table 1. Hence, we obtain the profile seen in Fig. 4, which is consistent with that obtained from the ELLPROF task in BBarolo. Thus, we can proceed to the Sersic modeling.
The Sersic profile is given by:
where and is the intensity at effective radius and n is the Sersic index.
@@
We fit a single-component Sersic model, and a two-component model to the derived profile and compare them with the Akaike information criterion (AIC) and Baysian information criterion (BIC) as well (Burnham et al. 2010). Both methods provide us with a metric of comparison for fit models. However, BIC heavily penalizes large number of parameters, and hence can provide a counter against overfitting and is great to test out a two-component fit.
@@
While comparing models, lower AIC and BIC values are preferred. AIC for the single component Sersic fit is 189.11 while for the two-component fit we have 57.42; AIC overwhelmingly supports the two-component fit. The BIC value for the single component fit is 28.65 and 22.49 for the two-component fit; in terms of probabilities given by:
@@
The two-component model is favored by a factor of 20 by BIC. Hence both criteria point to a two-component distribution of HI in WLM. Both in stars and gas, there are an inner compact component and an extended, more disk-like, outer component. The Sersic indices for the respective components are also in agreement as shown in the legend of Fig. 4 indicating that the two-components in both gas and stars are very similar.
@@
In Section 4 we model the galaxy using the tilted ring method, however the model is applied separately for the approaching and receding sides owing to the galaxy’s kinematic asymmetry.
@@ 4.1 The asymmetric Rotation Curve of WLM
Kepley et al. (2007) identified that WLM shows an asymmetric rotation curve on the approaching and receding sides. We can visually inspect the moment-1 map in Fig. 2 to see that the blue and the red side velocities show a symmetric rapid increase in the center, but the blue side ones taper down with sparsely packed contours while the red side keeps increasing rapidly.
The velocity contours in the moment-1 map significantly deviate from perpendicularity to the major axis, particularly in the central region. This, together with the fact that blue and red sides show different directions in the center could be a signature of a dense bar-like structure at the center, and we discuss this later in Section 6.
@@
Kepley et al. (2007) used a 2D tilted ring model to obtain the rotation curve on both sides of the galaxy separately. Within the first kiloparsec, the approaching side of the rotation curve rises above the receding side by km s-1, then the velocities cross over where the receding side rapidly rises like a solid body until the outskirts, while the approaching side shows a gentle rise. An asymmetric distribution of velocities is also apparent in the global velocity profile shown in Fig. 5, it points to a multi-component kinematic structure of the galaxy. This will be discussed in detail in section 5.
4.2 The expected impact of ram pressure to the asymmetric Rotation Curve
Yang et al. (2022) used the proper motion of WLM obtained by Battaglia et al. (2022) from Gaia DR3. The stripped gas is exactly in the opposite direction of the proper motion, which indicates that the ram pressure is applied in the same direction. As WLM moves through the IGM, its approaching side’s motion, which contributes to the rotation curve, is aligned against the direction of ram pressure, while the receding side’s motion is aligned along the direction of ram pressure. This scenario is depicted in Fig. 6.
@@ 4.3 The final Rotation Curve of WLM
To model the complex kinematics of WLM we adopt an iterative prescription based on the 3D tilted ring fitting tool TiRIFiC (Józsa et al. 2007). It allows for comprehensive disk modelling with a vast suite of available parameters to fit, with the possibility of granular control over the fitting procedure. We model the two halves of the disk separately owing to the asymmetry along the approaching and receding sides. All parameters are independently determined for both galaxy sides, which however share a common systemic velocity and center. The initial guesses for all parameters were used from Ianjamasimanana et al. (2020) and Kepley et al. (2007).
@@
In the first iteration, we provide initial guesses for rotational velocity, dispersion, surface brightness, systemic velocity, position angle, inclination, dynamical center and thickness for 14 rings on both sides of the galaxy. Only the surface brightness, systemic velocity and dynamical center are allowed to vary and to arrive at a best fit in the first iteration.
@@
Next, in the second iteration, we use the previously determined parameters as initial conditions, and let all of the mentioned parameters from the first iteration vary and then be determined further. Letting both the position and inclination vary all the way from the inner region to the outskirts should let us capture twist-like features in the galaxy. Then, to model non-circular motions, we fit for radial velocity and some second-order harmonics for velocity (RO2A, RO2P, RA2A, and RA2P in the TiRiFiC documentation) used by Spekkens and Sellwood (2007) to implement effects of bar streaming. Velocity harmonics were also used by Ianjamasimanana et al. (2020) to successfully model non-circular motions in WLM. We then compare the PV diagram and global velocity profile of the model for consistency, and then proceed to double the number of rings to 28 to capture smaller variations and rerun the fit with all the previous fit parameters as initial conditions. The model is then processed through TRMerrors222github.com/PeterKamphuis/TRM_errors to obtain errors. Acquiring errors on all the parameters for every ring is very computationally expensive. Hence, we compute errors only on rotational velocity, dispersion, inclination and position angle, presented in Appendix A. Values for inclination and position angle presented in table 1 are average values of all the fits presented in Appendix A
@@
Fig. 5 includes a comparison of the observed global velocity profile of WLM and of the TiRiFiC model obtained from the above procedure. The observation and the model seem to agree well. Comparing the PV diagrams of the model and the observation in Appendix B , we see that all observed features are recovered by the model, including the asymmetry between the approaching and receding sides as well as the 2 component-like split on the approaching side. We also provide moment maps and residuals for the TiRiFiC model in Appendix C.
@@
Kepley et al. (2007) used a 2D method to obtain their rotation curves, and did not fit for dispersion and thickness. A similar prescription using TiRiFiC and velocity harmonics has been used to model WLM earlier by Ianjamasimanana et al. (2020), but they model both sides together with a single disk.
@@
In Sections 5 and 6, we discuss in detail about the inner component of WLM and the possibility of a bar-like structure in the center. Such a structure can provide inner disk velocity asymmetries, and contribute to non-circular motions, which we have considered and fit for in our TiRiFiC modelling. Sofue and Kohno (2025) report that for the rotation curve of the Milky Way along its inner bar region, the variation along two different regions of the Galaxy is 10-15%. If a similar impact were to be assumed for WLM, it would be 2-3 km s-1 at 1.5 kiloparsec for our rotation curve. The asymmetry beyond 1.5 kiloparsec that has been found by Kepley et al. (2007) is consistent with our results.
@@
Beyond about 1.5 kiloparsec from the galaxy center, the restoring force is not strong enough from the galaxy center. Then, the approaching side loses velocity, and the receding side gains velocity as seen in the rotation curve’s divergence shown in Fig. 7, and it confirms that ram pressure may explain the asymmetric RC of WLM (see Fig. 6). We observe an increased dispersion on the eastern edge of the galaxy, which can be seen in the moment-2 map of Fig. 2. This region is also visible in the residual moment 1 and 2 maps presented in Appendix C. This is also the side of the galaxy on which ram pressure is exerted, complex turbulent gas motions and ram pressure induced star formation may explain this increased dispersion.
@@
When the rotation curve of a dwarf system does not exceed a few tens of km s-1, it becomes essential to apply the asymmetric-drift correction. More specifically, for gas discs with ratios of rotational velocity to velocity dispersion as low as Vrot/ 4, pressure support can significantly affect the kinematics and must be taken into account using the so-called asymmetric-drift correction (Binney and Tremaine 2011; Bureau and Carignan 2002; Oh et al. 2011). We implement the procedure described by Oh et al. (2015). However, this correction is only necessary to estimate the dynamical mass of the galaxy under the assumption of equilibrium, which will be examined in a future paper. Corrected rotation curves are shown in Fig. 7, uncorrected ones are presented in Appendix A.
4.4 The shift of the gas dynamical center due to ram pressure
@@
As described earlier, we let TiRiFiC find the dynamical center and systemic velocity in our iterative fitting procedure, common for all rings on either side of the galaxy, and it appears to be shifted by about 90 parsec from the stellar center in the direction of the ram pressure being applied to the galaxy, as reported in table 1. Further, we would like to note that the shift is within 1 beamsize and could be impacted by smearing.
@@
Ram pressure is expected to more affect the dynamics of the low column density gas-disk outskirts, while gas in the inner region of the galaxy is denser and has experienced more orbits. For instance, looking at the PV diagram in Fig. 8 we can see that the rotational velocity is roughly 20 km s-1 at 1 kpc, and we can estimate the number of orbits the gas has had at that radius to be 6.6 in 2 Gyr, while taking a rough velocity of 30 km s-1 at 2.5 kpc we get 4 orbits in 2 Gyr. Therefore, we expect the inner region to be more at dynamical equilibrium than the outer one, and we do confirm that the approaching and receding sides coincide in the inner region (see a discussion of galactic disk dynamics in Hammer et al. 2025). It is not uncommon for dwarf irregulars to have their HI centers slightly different from their optical centers derived from photometry (Patra et al. 2016). WLM seems substantially perturbed by ram pressure, with having lost 13% of its HI mass, just in the MeerKAT16 field of view, and this interaction may contribute to the shift of the dynamical center.
@@ 5 Kinematic Two-Component Decomposition
The global flux profile we present in Fig. 5 shows a strong peak at 10 km s-1 and a plateau-like feature between 0 and -25 km s-1. This suggests that there could be more than one rotating components in the gas body of WLM. Also, in the PV diagram presented in Fig.8, on the approaching side we can see that the flux distribution also suggests the presence of two distinct kinematic branches. Further evidence for more than one component could be seen if we observe the spectrum taken in the region beyond 1 kpc as seen in Fig. 8, which shows two distinct peaks, and indeed a two-component Gaussian fits much better to the spectrum in this region. On the receding side (Appendix D, row number 3), the second component may not be resolved, but the spectrum here is much broader than it is at the center of the PV diagram and is also better modeled by two components.
@@
To systematically decompose our data cube into these two distinct components, we have applied the above method to every pixel in the data cube instead of just the central slice of the PV diagram. First, the original data cube has been re-binned by a factor 4:1 to increase the signal-to-noise ratio. We then extract the spectrum at every pixel in the data cube and fit two Gaussian profiles to the spectrum in each pixel in the data cube akin to the fit seen in Fig. 8 (see also Appendix D).
@@
We segregate the components in two separate data cubes; at each pixel, we sort the component with the lower velocity of the two with respect to the systemic velocity into what we refer to as Component-1 and the larger velocity into what is referred to as Component-2. Beamsize smoothening kernel is applied to compensate for spurious fits from pixel to pixel. Then, we generate moment maps for these two components (see Fig. 9).
@@
The reproduction of the datacube after this two-component decomposition can be seen in the second row of Fig. 10. We can see that it near perfectly reproduces the column density distribution of HI in the moment-0 map, as well as the velocity field in the moment-1 map. There are minor local variations from the observations in the moment-2 dispersion map, but the global features are well reproduced. We also compute rotation curves for each component using the methodology already used for the global rotation curve (see Section 4 and Fig. 11).
@@
In Section 3, we obtained a 2 component Sersic fit on the total HI profile of WLM. This let us establish that the two components present in the stellar profile of WLM reported in Higgs et al. (2021) can be mapped onto the HI profile. Due to the high inclination of the disk, these components overlap in the global moment 0 map of WLM. The kinematic decomposition of the data lets us separate both the components cleanly, and obtain moment maps presented in Fig. 9. We can now fit a single Sersic profile individually to the two kinematically separated components and compare them to the Sersic fits from Section 3. This comparison can be seen in Fig. 12. Both the kinematic decomposition and morphological decomposition show a compact inner component and a sparse outer component. Though in both cases the distinction between a compact and a sparse component is very clear, the Sersic parameters do not match perfectly, which is expected due to projection effects and the overlap between the two components. We discuss the properties of the two components in detail in Section 6.
@@
We also obtained a global velocity profile for both components that we present in Fig. 13. Component-1 spans a smaller velocity range but contributes more to the sharp peak in the profile. This can be verified from the rotation curve, the velocity amplitude of this component is lower at all radii than that of Component-2. The latter is much more widely distributed in the velocity space, it extends beyond Component-1, and shows the differential rotation shape expected for a rotating disk which is the double horned profile. Fig. 13 shows that Component-2 has a slight velocity excess of few km s-1 on the receding side at about 40 km s-1, which might be due to ram pressure.

### v3

@@ 3 HI distribution in WLM
A visual inspection of our high-resolution moment-0 map in Fig. 2 shows us that WLM has a dense central structure that tapers into the rest of the gaseous disc. This is also seen in the corresponding position-velocity (PV) diagram in Fig. 3, where the dense gas core is seen along the steep velocity gradient at the centre of the galaxy.
@@
Higgs et al. (2021) found that the stellar radial profile of WLM in their g- and i-band survey is better fitted by 2 distinct morphological components instead of a single one, both components being described by a compact Sérsic profile in the inner region and a shallower Sérsic profile in the outer region. They tested whether a single component describes the profile better than two components by using the Akaike information criteria (AIC) and the Baysian Information Criteria (BIC). In their study for WLM a two-component model is almost 600 times more likely than the single component model, significantly favouring the two-component case. We repeated this test for the 1D gas profile of WLM.
@@
Classical ellipse fitting to obtain a 1D HI flux profile may not work well for a galaxy at a very high inclination such as WLM. Therefore, we used two separate methods to extract the 1D profile; We performed the ELLPROF task in BBarolo (Teodoro and Fraternali 2015), which uses standard tilted-ring models, and we also implemented the following simple procedure to get a profile to test for consistency.
@@
The semi-major axis of the galaxy was determined by measuring the flux in the moment-0 map along a straight line along the position angle (PA) reported in Table 1. The line was terminated when the flux reaches four times the RMS noise in the moment map. The RMS noise in the map was determined by the noise in one channel multiplied by the square root of the number of contributing channels (Ianjamasimanana et al. 2018).
@@
The semi-minor axis was calculated with the same procedure repeated along a line perpendicular to the PA along the centre of the galaxy and thus an axis ratio is calculated. A central pixel, corresponding to the chosen dynamical centre derived in Section 4.3 was chosen, and ellipses were placed on the image, their size starting with the axes lengths and then decreasing with a set value, while maintaining the axis ratio and PA of each. The image was broken into annular regions, with each ellipse being an annular boundary; the mean flux was calculated for each region, as was the standard error on the mean value. Note that the ellipses are not being fit, they are determining spatial bin regions.
@@
In Eq.1, SE, and n are the standard error, the standard deviation of the flux values and the number of pixels within each in the annulus, respectively.
@@
Thus the total uncertainty was calculated for each annulus using Eq. 2, with the RMS being calculated as described earlier in this section:
@@
These values are then corrected for inclination presented in Table 1. Hence, we obtain the profile seen in Fig. 4, which is consistent with that obtained from the ELLPROF task in BBarolo. Thus, we could proceed to the Sérsic modelling.
The Sérsic profile is given by:
where and is the intensity at effective radius and n is the Sérsic index.
@@
We fitted a single-component Sérsic model and a two-component model to the derived profile and compared them with the AIC and BIC (Burnham et al. 2010). Both methods provide us with a metric of comparison for fit models. However, the BIC heavily penalises large number of parameters, and hence can provide a counter against overfitting and is great to test out a two-component fit.
@@
While comparing models, lower AIC and BIC values are preferred. AIC for the single component Sérsic fit is 189.11 while for the two-component fit we have 57.42; AIC overwhelmingly supports the two-component fit. The BIC value for the single component fit is 28.65 and 22.49 for the two-component fit; in terms of probabilities given by:
@@
The two-component model is favoured by a factor of 20 by the BIC. Hence both criteria point to a two-component distribution of HI in WLM. Both in stars and gas, there are an inner compact component and an extended, more disc-like, outer component. The Sérsic indices for the respective components are also in agreement as shown in the legend of Fig. 4 indicating that the two components in the gas and stars are very similar.
@@
In Section 4 we model the galaxy using the tilted ring method. However, the model is applied separately for the approaching and receding sides owing to the galaxy’s kinematic asymmetry.
@@ 4.1 The asymmetric rotation curve of WLM
Kepley et al. (2007) identified that WLM shows an asymmetric rotation curve on the approaching and receding sides. We can visually inspect the moment-1 map in Fig. 2 to see that the blue and the red side velocities show a symmetric rapid increase in the centre, but the blue side ones taper down with sparsely packed contours while the red side keeps increasing rapidly.
The velocity contours in the moment-1 map significantly deviate from perpendicularity to the major axis, particularly in the central region. This, together with the fact that blue and red sides show different directions in the centre could be a signature of a dense bar-like structure at the centre, and we discuss this later in Section 6.
@@
Kepley et al. (2007) used a 2D tilted ring model to obtain the rotation curve on both sides of the galaxy separately. Within the first kiloparsec, the approaching side of the rotation curve rises above the receding side by km s-1, then the velocities cross over where the receding side rapidly rises like a solid body till the outskirts, while the approaching side shows a gentle rise. An asymmetric distribution of velocities is also apparent in the global velocity profile shown in Fig. 5, it points to a multi-component kinematic structure of the galaxy. This will be discussed in detail in section 5.
4.2 The expected impact of ram pressure on the asymmetric rotation curve
Yang et al. (2022) used the proper motion of WLM obtained by Battaglia et al. (2022) from Gaia DR3. The stripped gas is exactly in the opposite direction of the proper motion, which indicates that the ram pressure is applied in the same direction. As WLM moves through the IGM, its approaching side’s motion,is aligned against the direction of ram pressure, while the receding side’s motion is aligned along the direction of ram pressure. This scenario is depicted in Fig. 6.
@@ 4.3 The final rotation curve of WLM
To model the complex kinematics of WLM, we adopted an iterative prescription based on the 3D tilted ring fitting tool TiRIFiC (Józsa et al. 2007). It allows for comprehensive disc modelling with a vast suite of available parameters to fit, with the possibility of granular control over the fitting procedure. We modelled the two halves of the disc separately owing to the asymmetry along the approaching and receding sides. All parameters were independently determined for both galaxy sides, which however share a common systemic velocity and centre. The initial guesses for all parameters were used from Ianjamasimanana et al. (2020) and Kepley et al. (2007).
@@
In the first iteration, we provided initial guesses for rotational velocity, dispersion, surface brightness, systemic velocity, PA, inclination, dynamical centre and thickness for 14 rings on both sides of the galaxy. Only the surface brightness, systemic velocity and dynamical centre are allowed to vary and to arrive at a best fit in the first iteration.
@@
Next, in the second iteration, we used the previously determined parameters as initial conditions and let all of the mentioned parameters from the first iteration vary and then be determined further. Letting both the position and inclination vary all the way from the inner region to the outskirts should let us capture twist-like features in the galaxy. Then, to model non-circular motions, we fitted for radial velocity and some second-order harmonics for velocity (RO2A, RO2P, RA2A, and RA2P in the TiRIFiC documentation) used by Spekkens and Sellwood (2007) to implement effects of bar streaming. Velocity harmonics were also used by Ianjamasimanana et al. (2020) to successfully model non-circular motions in WLM. We then compared the PV diagram and global velocity profile of the model for consistency, and then doubled the number of rings to 28 to capture smaller variations and rerun the fit with all the previous fit parameters as initial conditions. The model was then processed through TRMerrors222github.com/PeterKamphuis/TRM_errors to obtain errors. Acquiring errors on all the parameters for every ring is very computationally expensive. Hence, we computed errors only on rotational velocity, dispersion, inclination and PA (see Appendix A). Values for inclination and PA presented in Table 1 are average values of all the fits presented in Appendix A
@@
Fig. 5 includes a comparison of the observed global velocity profile of WLM and of the TiRIFiC model obtained from the above procedure. The observation and the model seem to agree well. Comparing the PV diagrams of the model and the observation in Appendix B , we see that all observed features are recovered by the model, including the asymmetry between the approaching and receding sides as well as the two-component-like split on the approaching side. We also provide moment maps and residuals for the TiRIFiC model in Appendix C.
@@
Kepley et al. (2007) used a 2D method to obtain their rotation curves, and did not fit for dispersion and thickness. A similar prescription using TiRIFiC and velocity harmonics has been used to model WLM earlier by Ianjamasimanana et al. (2020), but they model both sides together with a single disc.
@@
In Sections 5 and 6 we discuss in detail the inner component of WLM and the possibility of a bar-like structure in the centre. Such a structure can provide inner disc velocity asymmetries, and contribute to non-circular motions, which we have considered and fit for in our TiRIFiC modelling. Sofue and Kohno (2025) report that for the rotation curve of the Milky Way along its inner bar region, the variation along two different regions of the Galaxy is 10-15%. If a similar impact were to be assumed for WLM, it would be 2-3 km s-1 at 1.5 kiloparsec for our rotation curve. The asymmetry beyond 1.5 kiloparsec that has been found by Kepley et al. (2007) is consistent with our results.
@@
Beyond about 1.5 kiloparsec from the galaxy centre, the restoring force is not strong enough from the galaxy centre. Then, the approaching side loses velocity and the receding side gains velocity as seen in the rotation curve’s divergence shown in Fig. 7, and it provides further evidence that ram pressure may explain the asymmetric rotation curve of WLM (see Fig. 6). We observe an increased dispersion on the eastern edge of the galaxy, which can be seen in the moment-2 map of Fig. 2. This region is also visible in the residual moment-1 and -2 maps presented in Appendix C. This is also the side of the galaxy on which ram pressure is exerted, complex turbulent gas motions and ram-pressure-induced star formation may explain this increased dispersion.
@@
When the rotation curve of a dwarf system does not exceed a few tens of km s-1, it becomes essential to apply the asymmetric-drift correction. More specifically, for gas discs with ratios of rotational velocity to velocity dispersion as low as Vrot/ 4, pressure support can significantly affect the kinematics and must be taken into account using the so-called asymmetric-drift correction (Binney and Tremaine 2011; Bureau and Carignan 2002; Oh et al. 2011). We implemented the procedure described by Oh et al. (2015). However, this correction is only necessary to estimate the dynamical mass of the galaxy under the assumption of equilibrium, which will be examined in a future paper. Corrected rotation curves are shown in Fig. 7, uncorrected ones are presented in Appendix A.
4.4 The shift in the gas dynamical centre due to ram pressure
@@
As described earlier, we let TiRIFiC find the dynamical centre and systemic velocity in our iterative fitting procedure, common for all rings on either side of the galaxy, and it appears to be shifted by about 90 parsec from the stellar centre in the direction of the ram pressure being applied to the galaxy, as reported in table 1. Further, we would like to note that the shift is within 1 beamsize and could be impacted by smearing.
@@
Ram pressure is expected to affect the dynamics of the low-column-density gas-disk outskirts more, while gas in the inner region of the galaxy is denser and has experienced more orbits. For instance, looking at the PV diagram in Fig. 8 we can see that the rotational velocity is roughly 20 km s-1 at 1 kpc, and we can estimate the number of orbits the gas has had at that radius to be 6.6 in 2 Gyr, while taking a rough velocity of 30 km s-1 at 2.5 kpc we get 4 orbits in 2 Gyr. Therefore, we expect the inner region to be more at dynamical equilibrium than the outer one, and we do confirm that the approaching and receding sides coincide in the inner region (see a discussion of galactic disc dynamics in Hammer et al. 2025). It is not uncommon for dwarf irregulars to have their HI centres slightly different from their optical centres derived from photometry (Patra et al. 2016). WLM seems substantially perturbed by ram pressure, with having lost 13% of its HI mass, just in the MeerKAT-16 field of view, and this interaction may contribute to the shift of the dynamical centre.
@@ 5 Kinematic two-component decomposition
The global flux profile we present in Fig. 5 shows a strong peak at 10 km s-1 and a plateau-like feature between 0 and -25 km s-1. This suggests that there could be more than one rotating components in the gas body of WLM. Also, in the PV diagram presented in Fig.8, on the approaching side we can see that the flux distribution also suggests the presence of two distinct kinematic branches. Further evidence of more than one component could be seen if we observe the spectrum taken in the region beyond 1 kpc as seen in Fig. 8, which shows two distinct peaks, and indeed a two-component Gaussian fits much better to the spectrum in this region. On the receding side (Appendix D, row 3), the second component may not be resolved, but the spectrum here is much broader than it is at the centre of the PV diagram and is also better modelled by two components.
@@
To systematically decompose our data cube into these two distinct components, we have applied the above method to every pixel in the data cube instead of just the central slice of the PV diagram. First, the original data cube has been re-binned by a factor of 4:1 to increase the signal-to-noise ratio. We then extracted the spectrum at every pixel in the data cube and fitted two Gaussian profiles to the spectrum in each pixel in the data cube akin to the fit seen in Fig. 8 (see also Appendix D).
@@
We segregated the components into two separate data cubes; at each pixel, we sorted the component with the lower velocity of the two with respect to the systemic velocity into what we refer to as Component-1 and the larger velocity into what is referred to as Component-2. A beamsize smoothing kernel is applied to compensate for spurious fits from pixel to pixel. Then, we generated moment maps for these two components (see Fig. 9).
@@
The reproduction of the datacube after this two-component decomposition can be seen in the second row of Fig. 10. We can see that it near perfectly reproduces the column density distribution of HI in the moment-0 map, as well as the velocity field in the moment-1 map. There are minor local variations from the observations in the moment-2 dispersion map, but the global features are well reproduced. We also computed rotation curves for each component using the methodology we used for the global rotation curve (see Section 4 and Fig. 11).
@@
In Section 3, we obtained a two-component Sérsic fit on the total HI profile of WLM. This let us establish that the two components present in the stellar profile of WLM reported in Higgs et al. (2021) can be mapped onto the HI profile. Due to the high inclination of the disc, these components overlap in the global moment-0 map of WLM. The kinematic decomposition of the data let us separate the two components cleanly and obtain the moment maps presented in Fig. 9. We could then fit a single Sérsic profile individually to the two kinematically separated components and compare them to the Sérsic fits from Section 3. This comparison can be seen in Fig. 12. Both the kinematic decomposition and morphological decomposition show a compact inner component and a sparse outer component. Though in both cases the distinction between a compact and a sparse component is very clear, the Sérsic parameters do not match perfectly, which is expected due to projection effects and the overlap between the two components. We discuss the properties of the two components in detail in Section 6.
@@
We also obtained a global velocity profile for the two components that we present in Fig. 13. Component-1 spans a smaller velocity range but contributes more to the sharp peak in the profile. This can be verified from the rotation curve, the velocity amplitude of this component is lower at all radii than that of Component-2. Component-2 is much more widely distributed in the velocity space, extends beyond Component-1, and shows the differential rotation shape expected for a rotating disc, i.e. a double-horn profile. Fig. 13 shows that Component-2 has a slight velocity excess of few km s-1 on the receding side at about 40 km s-1, which might be due to ram pressure.